(c) Find a particular solution curve Xp for the linear system cOs t X+ 2 sin t
Answer #1

Soluton Cniven cOslt) -2 X= + 28in(d) 4 Cos(t) and g a ain (t)J tet A- Firat we need to he matix A find Hu eigen value det (we take u,t thun u-t $CJ-N Complinetary So, u- thoefore. soluhion Ls Get fandomental set of Such hat we thave He the mahix 2eSince et costolt e 8in ()-2Co8() 5 e-at ginttot = - e-2 (2sin)+ Cost) +C and (5in() ) - 3C08(t) C e-8t and -3t ns1t) dd = 10melhal variaton of farameter He tonticular Hence Bolution is. gCos(t) 10 gin(d) 10 Cos(t) Bin(d) - p/ala,-k Sin(t)t9Cos(t) 5